# The Junction of Immersive Analytics and Virtual Reconstructions – A Case Study on the Mausoleum of Emperor Maxentius

### v1

@@ Abstract.
Virtual Archaeology has significantly evolved over the last decades through advancements in data acquisition and representation by, e. g., improved data recording technologies and virtual reality devices. Immersive environments provide novel ways to present historical events or objects with high visual quality for both, the general public as well as researchers. Here, we examine how the emerging field of Immersive Analytics can contribute to enhancing the understanding and exploration of archaeological data and explore the junction of Virtual Archaeology and Immersive Analytics, utilising reconstructions of the mausoleum of the late Roman emperor Maxentius in Rome for argumentation. Based on our work, we advocate the value of combining historical and computer science expertise for virtual reconstructions and immersive environments to facilitate deeper understanding and interactive exploration of archaeological data.
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JOCCHjournalvolume: xjournalnumber: xarticle: xpublicationmonth: 1ccs: Applied computing Architecture (buildings)ccs: Human-centered computing Visualization application domainsccs: Applied computing Computers in other domainsccs: Human-centered computing Computer supported cooperative workccs: Applied computing Archaeology
1. Introduction
@@
Terms like Virtual Archaeology, Virtual Restoration, and Virtual Reconstruction date back to the 1990s, when researchers investigated the potential of virtualisation for work on cultural heritage (Reilly, 1990; Prescott, 1997). Throughout the years, Virtual Archaeology has since grown into an entire field of research and seen immense development in its technological capabilities in terms of data acquisition and data representation, as well as several definitions of charters which employ meaningful standards for visualisations and virtual reconstructions (Lopez-Menchero and Grande, 2011; Denard, 2012). According to the Seville Charter (Lopez-Menchero and Grande, 2011), Virtual Archaeology concerns itself with finding comprehensive ways of visualising cultural heritage objects.
@@
The visualisation of archaeological objects began in the 2D realm and has since increasingly seen an extension through 3D to stereoscopic 3D (s3D) applications (e. g., models in virtual reality), since 2D visualisations often fall short of capturing the complexity and spatial relationships inherent in archaeological data. Immersive technologies can typically be classified by the reality-virtuality continuum by Milgram and Kishino (Milgram and Kishino, 1994). Most commonly, immersive technologies are understood to be either virtual reality (VR) as an entirely virtual environment, augmented reality (AR) as an overlay of additional information into the real world, or mixed reality (MR) as a blend between the virtual and real world. By leveraging immersive technologies, researchers can generate detailed and realistic 3D experiences of not only the sites and artefacts but also reconstructed objects, background information, or a contextually appropriate environment. This leads to a more immersive and interactive exploration of the past.
@@
Rome, as a focal point of archaeological work in Europe, enjoys a high popularity for such work, leading to an increase in more extensive reconstruction visualisations. An example of such an immersive model representation is Rome Reborn, a virtual reality project that has been running since 1996 at the University of Indiana, which reconstructs Rome around 320 AD (Frischer and Massey, 2022, pp. 23–25). However, since such reconstructions target a broader audience, their representations are often suggestive. The transparency of their underlying archaeological data origin tends to be selective, and their authenticity and reliability are therefore limited. This is only natural because it can not be the goal of an immersive environment targeted at the broader public to discuss different opinions of scholarship regarding the reconstruction (Liverani, 2021, pp. 13–14).
@@
However, to support scientific debates, digital reconstructions can be used to develop innovative questions on the historical perception and sociopolitical embedding of monuments and objects, visualise different chronological phases of archaeological sites, and reflect on historical monuments’ construction history and building process. A clear indication of different phases, information sources or certainty of validity can, among others, take the form of colour mappings or opacity of the respective components integrated into the visualisation (Bartz et al., 2016; Demetrescu, 2018; Pietroni and Ferdani, 2021).
@@
To combine the potential of immersive environments and scientific debates, the relatively young field of Immersive Analytics (IA) offers plenty of advancements. IA uses engaging and embodied tools to support data understanding (Marriott et al., 2018), explores the potential of immersive environments for data analysis, and develops and evaluates designs of such environments tailored towards the use in specific application areas. Therefore, IA combines topics from visual analytics, computer graphics, human-computer interaction, and virtual reality. As a consequence, several areas that exhibit potential for archaeological research, including s3D, collaborative analysis, and multimodal representation and interaction, can be considered a part of it (Bienroth et al., 2022; Nowke et al., 2013). The investigation of which navigation or interaction methods are most appropriate for a specific context (Drogemuller et al., 2018), how effective VR is in teaching contexts (Sari et al., 2023), or the impact of different modalities like audio or haptics (Ruotolo et al., 2013; Venkatesan et al., 2023) are some of its core questions. The answers to those questions are as relevant for cultural heritage dissemination and research as they are for other applications. They support understanding of findings and to develop hypotheses, which in a reconstruction context means deciding on more meaningful interpretations of the original historical intent and use. In addition to enhancing data representation and analysis, IA can help communicate results in a more engaging and accessible way, like annotating specific objects or positions or recording workflows. Creating immersive environments that allow people to explore sites, artefacts, and reconstructed objects virtually, to add annotations and landmarks for idea communication, and to obtain more than mere visual impressions of cultural heritage objects, supports bridging the gap between academia and the public. Current scientific debates can be communicated by condensing the underlying information and providing a more relatable and captivating experience.
@@
Since immersive experiences are broadly utilised in the context of dissemination, like museums or teaching, as well as for research (Boboc et al., 2022; Rodriguez-Garcia et al., 2024), and evidence suggests that discourse in Virtual Archaeology is mainly driven by advancing technologies (Münster, 2019), we discuss perspectives of how the mentioned potentials of IA can be utilised. We discuss how Immersive Analytics methods can be used for Virtual Archaeology, specifically for virtual reconstructions, and utilise three reconstructions of the Maxentius Mausoleum in Rome for our argumentation (see Fig. 1). Those reconstructions are based on ongoing scientific debates on what the mausoleum might have looked like in its prime, and we provide those in addition to software to inspect them in VR. Thus, our contribution is a perspective on where those two research directions can be merged to further discourse and show, by example, some of the possible concepts and techniques.
Maxentius A                          Maxentius B                        Maxentius C
@@
Figure 1. This image shows the three proposed models of the Mausoleum of Maxentius we discuss in this work. The first model (Maxentius A) closely follows Rasch’s reconstruction proposal (Rasch, 1984). The second model (Maxentius B) is based on Rasch’s model but has a closed dome and seven windows in the clerestory modelled after the Tor de’ Schiavi (Johnson, 2009). The third model (Maxentius C) tries to merge the dome and the windows of Tor de’ Schiavi to the proportions of the Maxentius Mausoleum.
@@
Renderings of all three digital reconstruction alternatives we propose next to each other with a focus on the different roofs, which pose the main difference between those models.
2. Background
Given the interdisciplinary nature of the topic, this section will introduce key vocabulary that is contextually relevant. In the work of Lopez-Menchero et al. (Lopez-Menchero and Grande, 2011), virtual reconstruction is defined as using a virtual model to recover a building or object made by humans visually. Here, we differentiate between digital reconstructions, virtual reconstructions, and immersive reconstructions to indicate the different contexts. Digital reconstruction is the process of digitally reconstructing an object through digital (computational) tools. Virtual reconstruction is the representation of a cultural heritage object in a virtual space to explore and interact with it. It is an application of digital reconstruction. Immersive reconstruction is the representation of a cultural heritage object in an immersive environment, that is, an environment along Milgram’s reality-virtuality continuum. It can be virtual, augmented, or mixed and can include additional techniques, such as data physicalisation, to provide an immersive experience.
One term relevant for reconstructions is anastylosis as part of virtual restoration, which describes the process of using in-situ findings as a reference, where in-situ means that they are found at their original cultural heritage site or, more specifically, directly “where they belong”. Those findings are then placed at their original position in the reconstruction and utilised to infer further missing components. This workflow is thoroughly described and used as part of the Extended Matrix by Demetrescu et al. (Demetrescu, 2018), who use colour coding to represent the reliability between different components of a reconstruction from still existing in-situ components which reside at their original position to purely hypothetical parts.
As we will discuss later in this work, virtual and immersive reconstructions often consist not only of the cultural heritage object itself but also of an artificial environment which, to some degree, resembles the environment where the object was formerly to be found. This concept adheres to the idea of situated analytics, which puts data representations in relation to germane objects, places, and persons to improve understanding and decision-making (Marriott et al., 2018, p. 185). For example, it might overlay an existing object with analysis information about it and is, thus, tightly linked to AR applications. Embodiment relates to the interaction of a human with a system through touch, gestures, or voice as a form of physical affordance rather than point and click interaction as it is typical for desktop PCs (Marriott et al., 2018, p. 3, p. 101). That said, the concept of multi-modality is closely related, as it encompasses input and output addressing multiple senses, such as vision, hearing, touch, kinesthesis, and (less commonly) smell. The psychological immersion, also called engagement (Marriott et al., 2018, p. 10), is another important concept, as there is evidence that task performance and recall are boosted through heightened engagement (Makowski et al., 2017). The last concept to mention is collaborative immersive analytics. In IA, many problems addressed require more than one person to solve (Marriott et al., 2018, p. 221), leading to multiple people working in the same virtual environment. This includes the collaboration of multiple people from different physical locations and even at various points in time.
3. Immersive Analytics for Virtual Archaeology
The popularity of immersive technologies, such as AR and VR, for museums or expositions, teaching and even research in archaeological contexts has grown (Boboc et al., 2022; Rodriguez-Garcia et al., 2024). As we focus on reconstructions in this work, we consider digital reconstruction models as well as the data recordings, which lead to those being the data we work with and seek to understand. Consequently, the general intersection with IA is clear: We use current technologies to present our data, and we want to optimise the experience to further a user’s ability to understand the reconstructions.
To better understand data, some level of engagement is desirable, as there is evidence for an increased recall of data (Mark et al., 2003). This can be achieved by using immersive VR environments, which are popular in different outreach-related domains (Frischer and Massey, 2022; Schofield et al., 2018; Kyrlitsias et al., 2020). Such environments have proven to create a heightened sense of presence, as already argued by Kyrlitsias et al. (Kyrlitsias et al., 2020), which matches findings from IA. For example, Ragan et al. argue that the combination of s3D and head-tracking helps spatial judgment accuracy (Ragan et al., 2012). The analysis they conducted in their study looked into the different effects of stereovision, head-tracking and bodily movement by tasking their participants with context-related tasks, in which the participant’s error rate was almost halved when using stereovision and head-tracking. Nowadays, such analysis can easily be done for cultural heritage dissemination environments, as many modern head-mounted display devices have integrated telemetry hardware like eye-tracking, and all of them have either internal or external head-tracking. Museums, for instance, could collect data to see which points of interest in their exhibit attract visitors. In general, an investigation could be conducted on how people perceive any given environment and whether the scene is overloaded or achieves the desired recall of the presented information.
AR is also already seeing applications in virtual archaeology  (Tscheu and Buhalis, 2016). Such environments create a heightened sense of presence, which matches findings from IA that situated analytics supports reasoning (ElSayed et al., 2015) by embedding data into the physical world (Willett et al., 2017). Several papers on situated analytics cover pipelines for creating corresponding environments (Kalkofen et al., 2007; Willett et al., 2017; Fleck et al., 2023), leading to the conclusion that these pipelines can be used for virtual reconstruction models as well. One widespread use for AR is mapping digital information onto a physical referent, an approach described by the conceptual model by Willett et al. (Willett et al., 2017). Given such a physical referent is a heritage site or a part of it, reconstructed components can be acquired from an existing data set like digital twins, digital reconstructions or even just raw point clouds, and the actual physical object can be supplemented by a virtual representation visible in the AR system. For example, a proximity-based magical lense visualisation (Lee et al., 2023) could be used, where supplement information for an object is only shown in an embedded fashion given a certain proximity of the user to the object. Other known or inferred information can be overlayed apart from reconstructed components. This could be, for example, the approximate age of an illustration or an estimated purpose of an object, which could be calculated in the background, or the “Logical World”, as Willett et al.  call it. Another information that could be provided that way is the different types of archaeological evidence as described by Demetrescu et al. (Demetrescu, 2018), from actual residuals of the degraded objects to inferences from typologically similar objects or lore. Those types of evidence can be mapped to different levels of certainty, from well-preserved to purely hypothetical, and should be highlighted appropriately in reconstruction visualisations (Demetrescu, 2018; Schäfer, 2018; Pietroni and Ferdani, 2021).
While visualisation is essential, current immersive environment devices offer a wide variety of possibilities to utilise different modalities for input and output. One of which is haptics, which supports sense-making for priorly unseen data (Taher et al., 2017). Encoding information like material density or even the age of different objects into the intensity of haptic feedback are possible channels for additional information. Furthermore, data physicalisation, for instance, in the form of 3D prints, is already being used to provide a deeper understanding of existing information (Klein et al., 2022), which gives users the possibility of actually touching their data. Such physicalisations could, in turn, be used as a way of interacting with reconstructions. For example, an AR device could recognise the region a user is pointing at to provide information about it. Another multi-modal channel for in- and output is sound, and there are various ways of using sound to map data (sonification) (Madhyastha and Reed, 1995). Having access to realistic sound models can already help to interpret which purpose a specific room might have served (McBride, 2014). Since ambient sound can reduce distraction and improve presence (Kern and Ellermeier, 2020), the identification and application of contextually appropriate ambient sounds which are based on a user position (in the digital world) or activities, as well as attention control through specifically designed noises can be used to improve dissemination quality or aid validity assessment.
Further intriguing aspects of IA environments are the possibilities that arise from development kits such as Unity (Unity Technologies, 2023) or Unreal Engine (Epic Games, 2024), which come with working physics engines and multi-user capabilities. Through physics engines, concepts like lighting can be realistically simulated to allow for the investigation of their influence on reconstructions. Through multi-user capabilities, multiple people who do not necessarily have to be at the same place (distributed collaboration) can look at reconstructions and interact with them, which augments the understanding of users (Hill, 1982; Mark et al., 2003). By integrating user interfaces, either simple VR versions of desktop-like interfaces or more natural ones (Wigdor, 2011), users can furthermore be given the power to manipulate objects or the environment as a whole directly. Definitions and a more generalised approach towards interaction systems can be found in IA-related work (Büschel et al., 2018). Such interaction systems can be utilised and combined to reasonable manipulation methods of a cultural heritage object, such as interactively changing the lighting conditions, adding annotations or filters for components, or reassembling pieces of the reconstructions in a playful manner (Houska et al., 2024). With a selection of those considerations, we will further discuss the example of the Maxentius Mausoleum and how these considerations can be applied to an environment for it.
4. Maxentius Mausoleum
Figure 2. The current state of the Maxentius Mausoleum. Only the Krypta still remains. The front porch is partly missing and has partly been used to build a farmhouse.
[…]
A photo we took of the Maxentius Mausoleum in 2024. Visible are the remains of the crypt from the outside as well as the farmhouse.
To provide insight into the significance of our archaeological example, we briefly discuss its historical and typological embedding and connection to the reign of Maxentius. The latter is relevant, as after nearly a century, during which the Roman emperors had been virtually absent from Rome and spent most of their time campaigning at the frontiers of the empire, the short rule of Maxentius (r. 306-312) gave Rome again the status of a permanent imperial residence. Maxentius styled himself as the conservator of Rome (conservator urbis suae) and added several large buildings such as the Maxentius Basilica and the temple of Venus and Rome, thermal baths and an extension of the imperial palace on the Palatine Hill to the architectural ensemble of the city (Leppin and Ziemssen, 2007, pp. 12-13, 51, 74-82, 87-95).
On the city’s outskirts, at the third milestone of the Via Appia leading southwards, Maxentius built a residential complex, which encompassed a villa suburbana with an attached well-preserved circus and a monumental mausoleum directed towards the Via Appia. The mausoleum, which forms the core of the complex, is surrounded by a slightly offset portico of 99.50 by 85 meters, of which substantial remains (as high as 11 meters) can still be seen (Rasch, 1984, p. 14) (see Fig. 2). However, the preserved remains of the mausoleum give the modern visitor only a limited idea of its former importance and structure. Just the foundations of the original building remain, overbuilt partly by a farmhouse of the 18th century (Rasch, 1984, pp. 5–6). Inside the building, an ambulatory crypt around an overarching pillar with niches inside is well-preserved. The layout of the building includes a round podium with a diameter of 33.25 meters. Adjacent is a front porch partly overbuilt by a farmhouse with a length of 17.00 meters and a width of 22.68 meters. Covered by the farmhouse, a vaulted cellar of the front porch with a diameter of 9.10 by 16.33 meters remains. The still-existing crypt has a diameter of 23.54 meters, a central column 9.70 meters wide, and seven niches (Johnson, 2009, pp. 86–90).
@@ 4. Maxentius Mausoleum
The monumental mausoleum was strongly connected to Maxentius’ rule and likely almost completed when Maxentius was killed in 312 AD (Johnson, 2009, pp. 92–93). Its few remaining traces provide an excellent case study for the application of virtual archaeology. Typologically, the mausoleum can be attributed to a type of podium round building, which – according to the architectural historian Jürgen Rasch – ultimately derives from the Pantheon. From 1977 to 1979, Rasch conducted an architectural survey of the site of the Maxentius Mausoleum and proposed a drawn reconstruction of the building modelled on the Pantheon with its specific architectural components: A podium with a hexastyle front porch (pronaos), a domed circular building with seven niches, and a circular dome opening (oculus) for the second storey (Rasch, 1993, p. 84). However, this Pantheon version of the Maxentius Mausoleum has been questioned by Mark Johnson, referring to a temporally and typologically similar building – another late antique imperial mausoleum at the via Praenestina (the so-called Mausoleum of Tor de’ Schiavi) (Rasch, 1993). This monument is located, like the Maxentius Mausoleum, on the outskirts of Rome, close to an imperial villa suburbana. It follows the design of a podium round building featuring a slightly smaller, tetrastyle front porch with a domed vault without an oculus but four windows in the clerestory. These different perspectives, into which Maxentius’ Mausoleum can be classified, result in different conceivable models for its reconstruction, which will be discussed below.
5. Immersive Analytics for the Maxentius Mausoleum
Regarding the Maxentius Mausoleum, we chose Rasch’s work as a foundation for a first reconstruction (Maxentius A), which closely follows his suggested measurements. The second model (Maxentius B) assumes a closed dome following Johnson’s proposal (Johnson, 2009, pp. 91-92) and adds seven round windows in the clerestory with a diameter of 1.80 meters considering the proportions of Tor de’ Schiavi (Rasch, 1993, p. 28). The third model (Maxentius C) adapts the flatter dome of Tor de’ Schiavi (Rasch, 1993, p. 68) and the windows with a larger diameter.
Figure 3. Cross-faded version of Maxentius A. Therefore, the different parts of the building are visible, and a photogrammetric recording of the current state of the crypt is embedded into the reconstruction.
[…]
A transparent version of a digital reconstruction based on Rasch’s work is put on a photogrammetry-based digital twin of the current state of the mausoleum.

### v2

@@ Abstract
Virtual Archaeology has significantly evolved over the last decades through advancements in data acquisition and representation, for example, by improved data recording technologies and virtual reality devices. Immersive environments provide novel ways to present historical events or objects with high visual quality for both the general public and researchers. Here, we examine how the emerging field of Immersive Analytics can contribute to enhancing the understanding and exploration of archaeological data and explore the junction of Virtual Archaeology and Immersive Analytics. We discuss a selection of features already used by the community and look into how optimizing those can facilitate the discourse and impression of cultural heritage objects. As an argumentational basis, we introduce and utilize three digital reconstruction interpretations of the mausoleum of the late Roman emperor Maxentius in Rome, which are based on prior scientific work and a typological embedding. Based on our work, we advocate the value of combining historical and computer science expertise to optimize immersive environments for virtual reconstructions to facilitate deeper understanding and interactive exploration of archaeological data.
Index Terms:
virtual archaeology; cultural heritage; 3D reconstruction; immersive analytics; virtual reality; Maxentius mausoleum;
I Introduction
@@
Terms like Virtual Archaeology, Virtual Restoration, and Virtual Reconstruction date back to the 1990s, when researchers investigated the potential of virtualization for work on cultural heritage [1, 2]. Throughout the years, Virtual Archaeology has since grown into an entire field of research and seen strong development in its technological capabilities in terms of data acquisition and data representation, as well as definitions of charters that employ meaningful standards for visualizations and virtual reconstructions [3, 4]. According to the Seville Charter [3], Virtual Archaeology concerns itself with finding comprehensive ways of visualizing cultural heritage objects.
@@
The visualization of archaeological objects began in the 2D realm and has since increasingly seen an extension through 3D to stereoscopic 3D (s3D) applications (e. g., models in virtual reality) since 2D visualizations often fall short of capturing the complexity and spatial relationships inherent in archaeological data. Immersive technologies can typically be classified by the reality-virtuality continuum by Milgram and Kishino [5]. Most commonly, immersive technologies are understood to be either virtual reality (VR) as an entirely virtual environment, augmented reality (AR) as an overlay of additional information into the real world, or mixed reality (MR) as a blend between the virtual and real world. By leveraging immersive technologies, researchers can generate detailed and realistic 3D experiences of not only the sites and artifacts but also reconstructed objects, background information, or a contextually appropriate environment.
@@
Rome, as a focal point of archaeological work in Europe, enjoys high popularity for such work, leading to an increase in more extensive reconstruction visualizations. An example of such an immersive model representation is Rome Reborn, a virtual reality project that has been running since 1996 at the University of Indiana, which reconstructs Rome around 320 AD [6, pp. 23–25]. The transparency of the underlying archaeological data origin of such public audience-oriented applications tends to be selective, and their authenticity and reliability are therefore limited. This is only natural because it can not be the goal of an immersive environment targeted at the broader public to discuss different opinions of scholarship regarding the reconstruction [7, pp. 13–14].
@@
To support scientific debates, digital reconstructions are useful for developing innovative questions on the historical perception and sociopolitical embedding of monuments and objects, visualizing different chronological phases of archaeological sites, and reflecting on historical monuments’ construction history and building process. A clear indication of different phases, information sources, or certainty of validity can, among others, take the form of color mappings or opacity of the respective components integrated into the visualisation [8, 9, 10].
@@
To combine the potential of immersive environments and scientific debates, the relatively young field of Immersive Analytics (IA) offers plenty of advancements. IA uses engaging and embodied tools to support data understanding [11], explores the potential of immersive environments for data analysis, and develops and evaluates designs of such environments tailored towards the use in specific application areas. IA combines topics from visual analytics, computer graphics, human-computer interaction, and virtual reality. As a consequence, several areas that exhibit potential for archaeological research, including s3D, collaborative analysis, and multimodal representation and interaction, can be considered a part of it [12, 13]. The investigation of which navigation or interaction methods are most appropriate for a specific context [14], how effective VR is in teaching contexts [15], or the impact of different modalities like audio or haptics [16, 17] are some of its core questions.
In addition to enhancing data representation and analysis, IA can help communicate results in a more engaging and accessible way, such as by investigating how to annotate specific objects or positions or how to collaborate in which contexts. Creating immersive environments that allow people to explore sites, artifacts, and reconstructed objects virtually, to add annotations and landmarks for idea communication, and to obtain more than mere visual impressions of cultural heritage objects supports bridging the gap between academia and the public. Thus, it is crucial to emphasize the optimization of those supporting features based on the cultural heritage item displayed and the user group towards which the virtual environment is targeted.
@@
Since immersive experiences are broadly utilized for dissemination, like museums or teaching, as well as for research [18, 19], and evidence suggests that discourse in Virtual Archaeology is mainly driven by advancing technologies [20], we discuss perspectives of how the mentioned potentials can be utilized. We discuss how Immersive Analytics methods can be used for Virtual Archaeology, specifically for virtual reconstructions. We use three reconstructions of the Maxentius Mausoleum in Rome for our argumentation (see Fig. 1). They are based on ongoing scientific debates about what the mausoleum might have looked like in its prime. We provide these reconstruction models in addition to software to inspect them in VR.
Thus, our contribution is comprised of a perspective on how Immersive Analytics research can help to assist archaeological discourse and the three digital reconstructions of the Maxentius mausoleum that we created. Furthermore, we provide an example software to demo some of the possible concepts and techniques.
@@
Figure 1: This image shows the three proposed models of the Mausoleum of Maxentius we discuss in this work. The first model (Maxentius A) closely follows Rasch’s reconstruction proposal [21]. The second model (Maxentius B) is based on Rasch’s model but has a closed dome and seven windows in the clerestory modelled after the Tor de’ Schiavi [22]. The third model (Maxentius C) tries to merge the dome and the windows of Tor de’ Schiavi to the proportions of the Maxentius Mausoleum.
II Background
The popularity of immersive technologies, such as AR and VR, for museums or expositions, teaching, and even research in archaeological contexts has grown [18, 19, 6, 23, 24]. As we focus on reconstructions in this work and its interdisciplinary nature, it’s important to clarify the different types of reconstructions in the digital realm. Digital reconstruction involves creating objects using computational tools, while virtual reconstruction is the use of virtual models to recover a building or man-made object visually [3]. Immersive reconstruction extends this further by embedding objects in environments along Milgram’s reality-virtuality continuum [5], such as AR or VR, to enhance user experience through interaction and context.
As is the nature of a reconstruction, it is not granted that all pieces of a monument, for instance, are known. Researchers have to complement the known parts with modeled ones, whose form and structure derive from different sources, like paintings, folklore, derivation from existing pieces, and so on. These different sources need to be pointed out, as Demetrescu describes in his work [9]. The sources of the reconstruction components, as well as the surrounding area and information on the object’s purpose or its material, are all data that can be transported via various means, especially so with the help of immersive technologies.
Immersive technologies can support reasoning [25] by embedding data into the physical world [26], the recall of data [27], or spatial judgment accuracy [28]. This is also relevant in Virtual Archaeology, as such environments have proven to create a heightened sense of presence in its’ context, as already shown by Kyrlitsias, Christofi, Michael-Grigoriou, Banakou, and Ioannou [29]. Presence, as a form of psychological immersion, can also be called engagement [30], and is an important concept, as there is evidence that task performance and recall are boosted through heightened engagement [31].
The research direction that concerns itself with investigating how to exploit those engaging tools to increase data understanding is called Immersive Analytics [11]. In their book on Immersive Analytics, Marriott et al. describe a number of relevant aspects apart from those already discussed. One of those is Situated analytics, which puts data representations in relation to germane objects, places, and persons to improve understanding and decision-making [32]. For example, it might overlay an existing object with analysis information about it and is, thus, tightly linked to AR applications. Another is the Embodiment, which relates to the interaction of a human with a system through touch, gestures, or voice as a form of physical affordance rather than point-and-click interaction as it is typical for desktop PCs [30]. That said, the concept of multi-modality is closely related, as it encompasses input and output addressing multiple senses, such as vision, hearing, touch, kinesthesis, and (less commonly) smell.
III Maxentius Mausoleum
Figure 2: The current state of the Maxentius Mausoleum. Only the Krypta still remains. The front porch is partly missing and has partly been used to build a farmhouse.
To provide insight into the significance of our archaeological example, we briefly discuss its historical and typological embedding and connection to the reign of Maxentius. The latter is relevant, as after nearly a century, during which the Roman emperors had been virtually absent from Rome and spent most of their time campaigning at the frontiers of the empire, the short rule of Maxentius (r. 306-312) gave Rome again the status of a permanent imperial residence. Maxentius styled himself as the conservator of Rome (conservator urbis suae) and added several large buildings such as the Maxentius Basilica and the temple of Venus and Rome, thermal baths, and an extension of the imperial palace on the Palatine Hill to the architectural ensemble of the city [33, pp. 12-13, 51, 74-82, 87-95].
@@ III Maxentius Mausoleum
The monumental mausoleum was strongly connected to Maxentius’ rule and likely almost completed when Maxentius was killed in 312 AD [22, pp. 92–93]. Its few remaining traces provide an excellent case study for the application of virtual archaeology. Typologically, the mausoleum can be attributed to a type of podium round building, which – according to the architectural historian Jürgen Rasch – ultimately derives from the Pantheon. From 1977 to 1979, Rasch conducted an architectural survey of the site of the Maxentius Mausoleum and proposed a drawn reconstruction of the building modeled on the Pantheon with its specific architectural components: A podium with a hexastyle front porch (pronaos), a domed circular building with seven niches, and a circular dome opening (oculus) for the second storey [34, p. 84]. However, this Pantheon version of the Maxentius Mausoleum has been questioned by Mark Johnson, referring to a temporally and typologically similar building – another late antique imperial mausoleum at the via Praenestina (the so-called Mausoleum of Tor de’ Schiavi) [34]. This monument is located, like the Maxentius Mausoleum, on the outskirts of Rome, close to an imperial villa suburbana. It follows the design of a podium round building featuring a slightly smaller, tetrastyle front porch with a domed vault without an oculus but four windows in the clerestory. These different perspectives, into which Maxentius’ Mausoleum can be classified, result in different conceivable models for its reconstruction, which will be discussed below.
IV Immersive Analytics for the Maxentius Mausoleum